OpenStudy (anonymous):

How would I factor : x^4-40x^2+144

OpenStudy (anonymous):

36+4=40 36*4=144

OpenStudy (anonymous):

I still don't understand what I do after I get the numbers 36 and 4

Directrix (directrix):

x^4 -40x^2 +144 = x^4 - 4x^2 - 36 x^2 + 144 = ( x^2 - 4 ) * ( x^2 - 36 ) = (x + 2)* (x - 2) * (x + 6 ) * ( x - ?) Fill in for the question mark. @abishop0018